OpenStudy (anonymous):

How do I convert 1.24 x 1024 molecules AgCH3COO (molecular weight 167 g/mol) to grams. will medal and fan!

OpenStudy (anonymous):

@JJHSstudent

OpenStudy (aaronq):

you wanna use \(n=\dfrac{N}{N_A}\) and \(n=\dfrac{m}{M}\) n=moles N=molecules \(N_A\)= avogadro's number m=mass M=molar mass

OpenStudy (anonymous):

I have n=2.08 what would be the next step? I got that from \[n=\frac{ N }{ N _{a} }\]

OpenStudy (aaronq):

so first find the moles, n (with that formula) then use the second to find the mass

OpenStudy (anonymous):

So I would do n*M=m?

OpenStudy (aaronq):

yes

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

just to check m=347.36?